Do It! Review 7-4b Monty Corporations management wants to maintain a minimum monthly cash balance of $9,600. At the beginning of September, the cash balance is $14,724, expected cash receipts for September are $116,640, and cash disbursements are expected to be $138,000 to be s138 wan How much cash, if any, must Manty borrow to maintain the desired minimum monthly balance? Determine your answer by using the basic form of the cash budget. MONTY CORPORATION Cash Budget
Answer #1
A B C D E F
2
3 Cash Budget can be prepared as follows:
4
5 Beginning Balance $14,724
6 Add: Expected Cash Receipt $116,640
7 Total cash available $131,364
8 Less: Cash Disbursement $138,000
9 Ending Cash Balance before loan adjustment ($6,636) =D7-D8
10 Add: Target Minimum Balance $9,600
11 Loan Borrowed $16,236 =MAX(D10-D9,0)
12

Formula sheet

A B C D E F
2
3 Cash Budget can be prepared as follows:
4
5 Beginning Balance 14724
6 Add: Expected Cash Receipt 116640
7 Total cash available =SUM(D5:D6)
8 Less: Cash Disbursement 138000
9 Ending Cash Balance before loan adjustment =D7-D8 =D7-D8
10 Add: Target Minimum Balance 9600
11 Loan Borrowed =MAX(D10-D9,0) =MAX(D10-D9,0)
12